OpenStudy (anonymous):

place the following fractions in order from least to greatest please: 1/7, 1/3, and 1/6

OpenStudy (anonymous):

1/7, 1/6, 1/3

Directrix (directrix):

1/7 = 6/42, 1/6 = 7/42, and 1/3 = 14/42. --> 1/7, 1/6, 1/3

OpenStudy (anonymous):

thanks medal for you and you!

Directrix (directrix):

... And, you, too.